A very broad elevator is going vertically up with constant acceleration of 2 m/s2 . At instant when its velocity is 4 m/sa ball is projected from the floor of the lift with a speed of 4 m/s relative to the floor at an elevation of 30 degree . The time taken by the ball to return the floor is( g = 10 m/s2 ) .
 
​a) ½ s b)1/3s c)1/4s d) 1s

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Net acceleration = g + a = > 12 m/s^2
Time of flight = 2usintheta /12 = > 2*4/sin30/12  = 1/3 seconds.